# Moving sum procedure for multiple change point detection in large factor models

### v1

@@ 3.3 Consistency in multiple change point estimation
Theorem 3 shows that the MOSUM procedure consistently estimates the total number and the locations of the change points. Here, we adopt a fixed, positive definite matrix $\widetilde{\mathbf{V}}$ in place of $\mathbf{V}_{k}$ which, without being a consistent estimator of the latter at some $k$, still leads to consistency in multiple change point detection. This flexibility in the choice of $\widetilde{\mathbf{V}}$ is particularly favourable since, as noted earlier in Section 2.2, the estimation of (time-varying) long-run covariance matrix for multivariate time series is challenging. At the same time, it incurs an additional multiplicative factor of $\omega_{T}^{(1)}$ in the detection lower bound placed on $d_{j}\sqrt{\gamma}$, see Assumption 7 (ii). Under a stronger moment assumption, we obtain the rate of estimation which is inversely proportional to the squared size of change as $|\widehat{k}_{j}-k_{j}|=O_{P}(d_{j}^{-2})$, This indicates that dominant changes are located with better accuracy, such as those accompanied by a change in the dimension of the factor space due to the introduction or disappearance of factors(s).
@@ 4.1 Tuning parameter selection
We set the threshold $D_{T,\gamma}$ as described in Theorem 3, namely $D_{T,\gamma}=\widetilde{D}_{T,\gamma}(\alpha)\cdot\omega_{T}^{(1)}$ where $\widetilde{D}_{T,\gamma}(\alpha)$ is given by (12) according to the asymptotic null distribution in Theorem 1. For $\omega_{T}^{(1)}$, we have considered $\log^{\kappa}(T/\gamma)$ with $\kappa\in\{0,0.1,0.2,0.3\}$, and observed that the choice of $\kappa=0.2$ returned stably good performance in all experiments. Due to the presence of this $\omega_{T}^{(1)}$, the selection of the significance level $\alpha\in(0,1)$ has relatively little influence and in all our studies, we set $\alpha=0.05$. Finally, for the detection rule in (4), we set $\eta=0.6$.
@@
Finally, in place of $\mathbf{V}_{k}$ in (3), we plug in the estimator $\widehat{\mathbf{V}}$ in (5) with bandwidth $m=\lfloor T^{1/4}\rfloor$. Due to the presence of change points, the number of pseudo factors $r$ can be large in which case inverting the $d\times d$-matrix $\widehat{\mathbf{V}}$ with $d=r(r+1)/2$, may bring numerical instability. Therefore, we explore two approaches, one performing the standardisation using the full matrix $\widehat{\mathbf{V}}$, and the other using its diagonal entries only, respectively referred to as MOSUM-full and MOSUM-diagonal in simulation studies. Note that MOSUM-diagonal meets the requirement in Theorem 3 provided that all diagonal entries of $\widehat{\mathbf{V}}$ are positive.
@@ 4.2 Settings
[…]
@@
    $\bm{\Sigma}_{1}=\bm{\Sigma}_{2}$ $=[\sigma_{1,ij}]\text{ \ with \ }\sigma_{1,ij}=\sigma_{1,ji}=\left\{\begin{array}[]{ll}0.9\sqrt{\sigma_{0,11}\sigma_{0,22}}&\text{for \ }(i,j)=(1,2),\\ 1.3^{2}\sigma_{0,55}&\text{for \ }(i,j)=(5,5),\\ 0.5^{|i-5|}\sqrt{\sigma_{0,ii}\sigma_{0,55}}&\text{for \ }1\leq i\leq 4,\\ \sigma_{0,ij}&\text{otherwise.}\end{array}\right.$
@@ 4.3 Results
Tables 14 report the summary of the results over $200$ realisations. Overall we observe that the MOSUM procedure (‘MOSUM’) performs competitively against BSCOV under (M1), and the former outperforms the latter in the other scenarios. In the presence of change points under (M2), we observe that BSCOV tends to over-estimate the number of change points particularly as $T$ increases, and this tendency becomes greater when the data are serially dependent (with $(\rho_{f},\rho_{e})=(0.7,0.3)$). In the absence of change points under (M3), BSCOV detects spurious change points when $T$ is smaller. On the contrary, MOSUM correctly estimates the number of change points on most realisations even when the data are serially dependent, regardless of whether $R=0$ or not. Additionally, we observe that $k_{j}$’s which are associated with changes in the number of pseudo factors (such as $k_{2}$ under (M1) and $k_{3}$ under (M2)) are estimated with higher accuracy, which agrees with the observations made in Duan et al., (2023) and also below Theorem 3.
Between MOSUM-full and MOSUM-diagonal, the latter demonstrates better accuracy in estimating both the total number and locations of change points when $R\geq 1$. This is explained by that $d=r(r+1)/2$ is as large as $d=28$ under (M1) and $d=21$ under (M2). Retaining the diagonal elements of $\widehat{\mathbf{V}}$ only, effectively performs standardisation without suffering from the numerical instability inherent in inverting a large matrix. When $R=0$, MOSUM-full performs marginally better as here, the factor number is kept at $r=3$ under (M3), leading to $d=6$.
Table 1: (M1) with $R=2$: Summary of change point estimators returned by MOSUM and BSCOV. The results for BSCOV have been taken from Li et al., (2023).
@@ 5 Real data application
The panel data is unbalanced with the dimension $N=72$ being considerably smaller than the sample size $T=4312$, a situation that does not favour the use of the sub-sampling-based factor number estimator as pointed out by Onatski, (2024). Indeed, it returns $r=7$ from the dataset which may be an over-estimate. An alternative estimator based on taking the ratio of successive eigenvalues of $(NT)^{-1}(\mathbf{X}\mathbf{X}^{\top})$ returns $r=1$ (Ahn and Horenstein,, 2013), which implies that any change point we detect is solely due to heteroscedasticity of the factor. Therefore, we apply the proposed MOSUM procedure with varying $r\in\{1,\ldots,7\}$ while setting other tuning parameters as described in Section 4.1; this gives a bandwidth $\gamma=227$ corresponding almost to one trading year.
Figure 1 illustrates the series of standardised MOSUM statistics $D_{T,\gamma}^{-1}\cdot\mathcal{T}_{N,T,\gamma}(k),\,\gamma\leq k\leq T-\gamma$, obtained for different values of $r$, where the standardisation is applied to ensure that the MOSUM series derived from $\textsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ of different dimensions are comparable. See also Table 5 for the list of change point estimators. We observe that with increasing $r$, the procedure detects more change points. Further, denoting by $\widehat{\mathcal{K}}(r)$ the set of change point estimators detected with the factor number $r$, generally it holds that $\widehat{\mathcal{K}}(r)\subset\widehat{\mathcal{K}}(r^{\prime})$ for any $r<r^{\prime}$, when accommodating the possible bias in the change point estimators (up to $3$ months). There are prominent estimators in mid-2008 and 2009 detected invariably with all $r$, which are associated with the Great Financial Crisis in 2007–2009. With $r\geq 5$, 2020-02-24 and 2021-01-19 emerge as change points, which are accounted for by the stock market crash in February 2020 and the ensuing recession due to the COVID-19 pandemic.
@@
Figure 1: US blue chip data: Standardised MOSUM statistics obtained with $r\in\{1,\ldots,7\}$ (top to bottom). Vertical lines denote the change point estimators and the horizontal line is at $y=1$.
@@
Table 5: US blue chip data: List of change point estimators obtained with $r\in\{1,\ldots,7\}$.
$r=1$ 2008-07-07, 2009-06-05
$r=2$ 2008-07-03, 2009-06-08
$r=3$ 2008-07-03, 2009-06-08
$r=4$ 2008-07-03, 2009-06-04, 2012-10-05
$r=5$ 2008-07-03, 2009-06-04, 2012-07-13, 2016-05-20, 2017-06-08, 2020-02-24, 2021-01-19
$r=6$ 2006-02-01, 2007-03-09, 2008-07-03, 2009-08-07, 2011-11-10, 2012-10-05, 2014-08-19
2016-05-20, 2017-06-08, 2020-02-24, 2021-01-19
$r=7$ 2006-02-01, 2007-03-09, 2008-09-12, 2009-05-28, 2011-11-10, 2012-10-05, 2014-08-19
2016-02-17, 2017-06-08, 2020-02-24, 2021-01-19
@@ 6 Conclusions
The paper proposes a MOSUM procedure for change point analysis under a static factor model that is popularly adopted in econometrics and statistics. In addition to deriving the asymptotic null distribution of the maximally selected MOSUM statistic, we establish the consistency of the procedure in multiple change point estimation with the accompanying rate of estimation, contributing to the relatively scarce literature on multiple change point detection in factor models. On a range of simulated datasets and in a real data application, we demonstrate the competitiveness of the proposal empirically. At the same time, the success of the proposed single-scale MOSUM procedure hinges on the availability of the bandwidth $\gamma$ that fulfils a set of assumptions, which may not exist in the presence of multiscale change points (Cho and Kirch,, 2024). One natural avenue for an extension is to apply the MOSUM procedure with a range of bandwidths, which we leave for future research.
@@ A.3 Proof of Proposition 2
Since $\mathbf{g}_{t}$ is an $\mathcal{L}_{8+\epsilon}$-decomposable Bernoulli shift with $a>2$, it is easy to see (cfr. the proof of Lemma A.9) that $\mathsf{Vech}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})$ is an $\mathcal{L}_{4+\epsilon/2}$-decomposable Bernoulli shift, also with $a>2$. The covariance summability of Bernoulli shifts (see e.g. Lemma D.4 in Horváth and Trapani,, 2023) entails that
@@
and again, we set $r=d=1$ for simplicity. Under Assumption 1 (i), it is easy to see that for all $j\geq 0$, the sequence $S_{t,\ell}=\widetilde{U}_{t}\widetilde{U}_{t-\ell}^{\top}-\mathsf{E}(\widetilde{U}_{t}\widetilde{U}_{t-\ell}^{\top})$ is an $\mathcal{L}_{2}$-decomposable Bernoulli shift with $a>2$. Hence
$\mathsf{E}\left(\left|\frac{1}{T}\sum_{\ell=1}^{m}\sum_{t=\ell+1}^{T}S_{t,\ell}\right|^{2}\right)\leq T^{-2}m\sum_{\ell=1}^{m}\mathsf{E}\left(\left|\sum_{t=\ell+1}^{T}S_{t,\ell}\right|^{2}\right)\leq\frac{c_{0}m^{2}}{T},$
@@
$\left|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right|=O_{P}\left(\frac{m}{T^{1/2}}\right).$ (A.55)
@@
$\left|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right|=O_{P}\left(\frac{m}{C_{NT}}\right)$ (A.56)
@@
$\left|\frac{1}{T}\sum_{t=1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}-\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t}^{\top}\right)\right|=O_{P}\left(\frac{1}{C_{NT}}\right).$ (A.57)
@@
$\left|\widehat{\mathbf{V}}-\mathbf{V}\right|=O_{P}\left(\frac{m}{C_{NT}}\right)+O\left(\frac{1}{m}\right)=O_{P}\left(\frac{1}{\log(T/\gamma)}\right),$

### v2

@@ 3.3 Consistency in multiple change point estimation
Theorem 3 shows that the MOSUM procedure consistently estimates the total number and the locations of the change points. Here, we adopt a fixed, positive definite matrix $\widetilde{\mathbf{V}}$ in place of $\mathbf{V}_{k}$ which, without being a consistent estimator of the latter at some $k$, still leads to consistency in multiple change point detection. This flexibility in the choice of $\widetilde{\mathbf{V}}$ is particularly favourable since, as noted earlier in Section 2.2, the estimation of (time-varying) long-run covariance matrix for multivariate time series is challenging. While the asymptotic null distribution in Theorem 1 allows for testing the null hypothesis of no change point with the family-wise error controlled, strengthening of the threshold is necessary for consistently detecting the number of change points, see e.g. Eichinger and Kirch, (2018) and Bai et al., (2024) where they set $\alpha=\alpha_{T}\to 0$ at a suitable rate. Instead, we introduce an additional multiplicative factor of $\omega_{T}^{(1)}\to\infty$ in the threshold $D_{T,\gamma}$; see Section 4.1 where we discuss the choice of the threshold. Under a stronger moment assumption, we obtain the rate of estimation which is inversely proportional to the squared size of change as $|\widehat{k}_{j}-k_{j}|=O_{P}(d_{j}^{-2})$, This indicates that dominant changes are located with better accuracy, such as those accompanied by a change in the dimension of the factor space due to the introduction or disappearance of factors(s).
@@ 4.1 Tuning parameter selection
We set the threshold $D_{T,\gamma}$ as described in Theorem 3, namely $D_{T,\gamma}=\widetilde{D}_{T,\gamma}(\alpha)\cdot\omega_{T}^{(1)}$ where $\widetilde{D}_{T,\gamma}(\alpha)$ is given by (12) according to the asymptotic null distribution in Theorem 1. As for $\omega_{T}^{(1)}$, we have considered $\log^{\kappa}(T/\gamma)$ with $\kappa\in\{0,0.1,0.2,0.3\}$, and observed that the choice of $\kappa=0.2$ returned stably good performance in all experiments, see Appendix B.2 for full detail. Compared to the choice of $\omega_{T}^{(1)}$, the selection of the fixed significance level $\alpha\in(0,1)$ has relatively little influence and in all our studies, we set $\alpha=0.05$. Finally, for the detection rule in (4), we set $\eta=0.6$.
@@
Finally, in place of $\mathbf{V}_{k}$ in (3), we plug in the estimator $\widehat{\mathbf{V}}$ in (5) with bandwidth $m=\lfloor T^{1/4}\rfloor$. Due to the presence of change points, the number of pseudo factors $r$ can be large in which case inverting the $d\times d$-matrix $\widehat{\mathbf{V}}$ with $d=r(r+1)/2$, may bring numerical instability, see Kirch et al., (2015) for the alternative approaches to handling similar difficulties in multivariate time series segmentation. Therefore, we explore two approaches, one performing the standardisation using the full matrix $\widehat{\mathbf{V}}$, and the other using its diagonal entries only, respectively referred to as MOSUM-full and MOSUM-diagonal. We remark that MOSUM-diagonal meets the requirement in Theorem 3 provided that all diagonal entries of $\widehat{\mathbf{V}}$ are positive. Our numerical experiments indicate that MOSUM-diagonal is to be preferred between the two, see Section 4.3 for further discussions.
@@ 4.2 Settings
    Adopted from Li et al., (2023), we fix $T=400$, $N=200$ and $r_{0}=5$, and introduce $R=2$ change points at $(k_{1},k_{2})=(133,267)$ as follows: $X_{it}=\chi_{it}+\sqrt{0.5}e_{it}$, where
@@
    $\bm{\Sigma}_{1}=\bm{\Sigma}_{2}$ $=[\sigma_{1,ij}]\text{ \ with \ }\sigma_{1,ij}=\sigma_{1,ji}=\left\{\begin{array}[]{ll}0.9\sqrt{\sigma_{0,11}\sigma_{0,22}}&\text{for \ }(i,j)=(1,2),\\ 1.3^{2}\sigma_{0,55}&\text{for \ }(i,j)=(5,5),\\ 0.5^{|i-5|}\sqrt{\sigma_{0,ii}\sigma_{0,55}}&\text{for \ }1\leq i\leq 4,\\ \sigma_{0,ij}&\text{otherwise,}\end{array}\right.$
@@ 4.3 Results
Figure 1: (M2) Histogram of the change point estimators returned by MOSUM-diagonal, BSCOV and BDH when $N=100$, $(\rho_{f},\rho_{e})=(0,0)$ and varying $T\in\{400,600,800,1000\}$ (top to bottom). The scaled locations of the true change points, $k_{j}/T$, at $(1/4,1/2,3/4)$ are marked by vertical dotted lines.
Figure 2: (M2) Histogram of the change point estimators returned by MOSUM-diagonal, BSCOV and BDH with $p=100$, $(\rho_{f},\rho_{e})=(0.7,0.3)$ and varying $T\in\{400,600,800,1000\}$ (top to bottom). The scaled locations of the true change points, $k_{j}/T$, at $(1/4,1/2,3/4)$ are marked by vertical dotted lines.
@@
  1. (i)
    Proposed by Li et al., (2023), BSCOV scans for changes in the covariance of $\{\mathbf{g}_{t}\}$ under (2) via an extension of the binary segmentation, produces a path of solutions and selects the final change point model by minimising an information criterion.
  2. (ii)
    Proposed by Bai et al., (2024), BDH recursively applies the likelihood ratio test via binary segmentation to detect the multiple change points under (2).
Both methods are applied with default tuning parameters and in-built factor number estimators that are based on the information criterion proposed by Bai and Ng, (2002); for BDH, we set the proportion of the data trimmed off at each recursion to be $0.1$. Tables 14 report the summary of the results over $200$ realisations, and Figures 12 plot the histograms of change point estimators returned by the proposed MOSUM procedure, BSCOV and BDH under (M2) when $N=100$; see also Appendix B for the additional results.
Overall we observe that the MOSUM procedure (‘MOSUM’) demonstrates competitive performance across all scenarios, both in detection and estimation. BSCOV tends to return spurious estimators under (M2) and (M3) when serial dependence is present with $(\rho_{f},\rho_{e})=(0.7,0.3)$. On the other hand, BDH suffers from lack of detection power against those changes that transform the loading matrix while do not alter the number of local factors, such as $k_{1}$ under (M1) and (M2), particularly for smaller $T$. For such a change point, MOSUM is able to detect its presence although with less accuracy. Generally, $k_{j}$’s which are associated with changes in the number of pseudo factors (such as $k_{2}$ under (M1) and $k_{3}$ under (M2)) are estimated with higher accuracy, which agrees with the observations made in Duan et al., (2023) and also below Theorem 3.
Between MOSUM-full and MOSUM-diagonal, the latter demonstrates better accuracy in estimating both the total number and locations of change points when $R\geq 1$. This is explained by that $d=r(r+1)/2$ is as large as $d=28$ under (M1) and $d=21$ under (M2). Retaining the diagonal elements of $\widehat{\mathbf{V}}$ only, effectively performs standardisation without suffering from the numerical instability inherent in inverting a large matrix. When $R=0$, MOSUM-full performs marginally better as here, the factor number is kept at $r=3$ under (M3), leading to $d=6$. Based on these observations, in practice, we recommend the use of MOSUM-diagonal when the number of (pseudo) factors is moderately large.
Table 1: (M1) with $R=2$: Summary of change point estimators returned by MOSUM, BSCOV and BDH. The results for BSCOV have been taken from Li et al., (2023).
@@ 5 Real data application
The sub-sampling-based factor number estimator discussed in Section 4.1 returns $r=7$ as the factor number. However, the panel data is unbalanced with the dimension $N=72$ being considerably smaller than the sample size $T=4312$, a situation that does not favour the sub-sampling approach as pointed out by Onatski, (2024). The information criteria of Bai and Ng, (2002) return $r=5$, while the approach based on inspecting the ratio of successive eigenvalues (Ahn and Horenstein,, 2013) returns $r=1$, which implies that any change point we detect would solely be attributed to heteroscedasticity of the single factor. While the former is known to detect weakly pervasive factors (Bai and Ng,, 2023), the latter tends to recover only strongly pervasive ones.
In the presence of some uncertainty in the number of factors, a situation commonly faced in real data analysis, we choose to apply the proposed MOSUM procedure with varying $r\in\{1,\ldots,7\}$ and inspect its outputs. We set other tuning parameters as described in Section 4.1 and adopt the standardisation based on the diagonal entries of $\widehat{\mathbf{V}}$ only (referred to as ‘MOSUM-diagonal’ in Section 4); this gives a bandwidth $\gamma=227$ corresponding almost to one trading year. Figure 3 illustrates the series of standardised MOSUM statistics $D_{T,\gamma}^{-1}\cdot\mathcal{T}_{N,T,\gamma}(k),\,\gamma\leq k\leq T-\gamma$, obtained for different values of $r$, where the standardisation is applied to ensure that the MOSUM series derived from $\textsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ of different dimensions are comparable. See also Figure 4 and Table 5 for the visualisation and the list of change point estimators returned by MOSUM and those competitors considered in the simulation studies, namely BSCOV (Li et al.,, 2023) and BDH (Bai et al.,, 2024).
It is noteworthy that the sets of change point estimators output by the MOSUM procedure are nested as the number of pseudo factors increases. That is, denoting by $\widehat{\mathcal{K}}(r)$ the set of change point estimators with $r$ as the factor number, we have $\widehat{\mathcal{K}}(r)\subset\widehat{\mathcal{K}}(r^{\prime})$ for any $r<r^{\prime}$, when accommodating the possible bias in the change point estimators (up to $3$ months). Specifically, with $r\leq 3$, we detect two prominent changes in mid-2008 and 2009 which, being associated with the Great Financial Crisis in 2007–2009, are detected invariably with all $r$. With $r=4$, we additionally detect 2012-10-05 as a change point, which is subsequently detected for all $r\geq 5$. With $r\geq 5$, 2020-02-24 and 2021-01-19 emerge as change points, which are accounted for by the stock market crash in February 2020 and the ensuing recession due to the COVID-19 pandemic. With $r\in\{6,7\}$, MOSUM outputs almost identical sets of change point estimators. These results offer an interpretation as to how different factors ‘encode’ different structural changes, and demonstrate that the MOSUM procedure is insensitive to the specified number of factors within certain ranges (i.e. $\{1,2,3\}$, $\{6,7\}$).
Similarly to MOSUM, BSCOV also returns nested sets of change point estimators as $r$ increases, and many of its estimators overlap with those returned by MOSUM. When $r=7$, the estimators returned by MOSUM form a subset of those returned by BSCOV, and some changes detected solely by BSCOV do not appear as estimators detected with $r<7$ by any method. BDH estimates the number of factors by the information criterion of Bai and Ng, (2002) at each iteration of the binary segmentation algorithm and as such, when applied with a fixed number of factors, its output lacks the nested property.
@@
Figure 3: US blue chip data: Standardised MOSUM statistics obtained with $r\in\{1,\ldots,7\}$ (top to bottom). Vertical lines denote the change point estimators returned by MOSUM and the horizontal line is at $y=1$.
Figure 4: US blue chip data: Volatilities from the $72$ companies. Vertical lines denote the change point estimators detected by MOSUM, BSCOV and BDH (top to bottom), with $r=7$.
@@
Table 5: US blue chip data: List of change point estimators obtained with $r\in\{1,\ldots,7\}$ by MOSUM, BSCOV (Li et al.,, 2023) and BDH (Bai et al.,, 2024).
$r$ MOSUM BSCOV BDH
$1$ 2008-07-07, 2009-06-05 2008-09-12, 2009-03-25
$2$ 2008-07-07, 2009-06-05 2008-09-12, 2009-03-25
$3$ 2008-07-07, 2009-06-05 2008-09-12, 2009-03-25 2008-06-27, 2029-08-07
$4$ 2008-07-03, 2009-06-04, 2012-10-05 2008-09-12, 2009-05-08, 2012-10-05 2008-07-03, 2009-07-23, 2012-06-21,
2014-09-29, 2020-02-24
$5$ 2008-07-03, 2009-06-04, 2012-07-13, 2008-09-12, 2009-05-08, 2012-10-08, 2007-10-31, 2012-10-05
2016-05-20, 2017-06-08, 2020-02-24, 2014-10-24, 2020-03-05
2021-01-19
$6$ 2006-02-01, 2007-03-09, 2008-07-03, 2007-01-11, 2008-09-12, 2009-05-08, 2012-10-05
2009-08-07, 2011-11-10, 2012-10-05, 2011-12-23, 2012-10-08, 2014-08-29,
2014-08-19, 2016-05-20, 2017-06-08, 2020-03-05
2020-02-24, 2021-01-19
$7$ 2006-02-01, 2007-03-09, 2008-09-12, 2006-05-18, 2007-02-22, 2008-09-12, 2012-09-26
2009-05-28, 2011-11-10, 2012-10-05, 2009-05-08, 2010-04-26, 2011-12-23,
2014-08-19, 2016-02-17, 2017-06-08, 2012-10-08, 2014-08-29, 2016-02-17,
2020-02-24, 2021-01-19 2017-08-28, 2019-10-10, 2020-03-05, 2021-04-01
@@ 6 Conclusions
This paper proposes a MOSUM procedure for change point analysis under a static factor model that is popularly adopted in econometrics and statistics. In addition to deriving the asymptotic null distribution of the maximally selected MOSUM statistic, we establish the consistency of the procedure in multiple change point estimation with the accompanying rate of estimation, contributing to the relatively scarce literature on multiple change point detection in factor models. On a range of simulated datasets and in a real data application, we demonstrate the competitiveness of the proposal empirically. At the same time, the success of the proposed single-scale MOSUM procedure hinges on the availability of the bandwidth $\gamma$ that fulfils a set of assumptions, which may not exist in the presence of multiscale change points (Cho and Kirch,, 2024). One natural avenue for an extension is to apply the MOSUM procedure with a range of bandwidths, which we leave for future research.
@@ A.3 Proof of Proposition 2
Since $\mathbf{g}_{t}$ is an $\mathcal{L}_{8+\epsilon}$-decomposable Bernoulli shift with $a>2$, it is easy to see (cfr. the proof of Lemma A.9) that $\mathsf{Vech}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})$ is an $\mathcal{L}_{4+\epsilon/2}$decomposable Bernoulli shift, also with $a>2$. The covariance summability of Bernoulli shifts (see e.g. Lemma D.4 in Horváth and Trapani,, 2023) entails that
@@
and again, we set $r=d=1$ for simplicity. Under Assumption 1 (i), it is easy to see that for all $j\geq 0$, the sequence $\mathbf{S}_{t,\ell}=\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}-\mathsf{E}(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top})$ is an $\mathcal{L}_{2}$-decomposable Bernoulli shift with $a>2$. Hence
$\mathsf{E}\left(\left\|\frac{1}{T}\sum_{\ell=1}^{m}\sum_{t=\ell+1}^{T}\mathbf{S}_{t,\ell}\right\|^{2}\right)\leq T^{-2}m\sum_{\ell=1}^{m}\mathsf{E}\left(\left\|\sum_{t=\ell+1}^{T}\mathbf{S}_{t,\ell}\right\|^{2}\right)\leq\frac{c_{0}m^{2}}{T},$
@@
$\left\|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right\|=O_{P}\left(\frac{m}{T^{1/2}}\right).$ (A.55)
@@
$\left\|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right\|=O_{P}\left(\frac{m}{C_{NT}}\right)$ (A.56)
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}-\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t}^{\top}\right)\right\|=O_{P}\left(\frac{1}{C_{NT}}\right).$ (A.57)
@@
$\left\|\widehat{\mathbf{V}}-\mathbf{V}\right\|=O_{P}\left(\frac{m}{C_{NT}}\right)+O\left(\frac{1}{m}\right)=O_{P}\left(\frac{1}{\log(T/\gamma)}\right),$